40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 26 17:24:24 2017, Steve, Update, General, power outage powerOutage.png
    Reply  Thu Dec 28 17:22:02 2017, gautam, Update, General, power outage - CDS recovery 28.png
       Reply  Thu Jan 4 14:39:50 2018, gautam, Update, General, power outage - timing error 
          Reply  Fri Jan 5 21:54:28 2018, rana, Update, General, power outage - timing error huh.png
             Reply  Fri Jan 5 22:19:53 2018, gautam, Update, General, power outage - timing error 
                Reply  Sat Jan 6 18:27:37 2018, gautam, Update, General, power outage - IFO recovery EX_QPD_Quad1_Faulty.pdfTrillium_faulty.pdf
                   Reply  Mon Apr 16 22:09:53 2018, Kevin, Update, General, power outage - BLRM recovery AA.jpgADC.jpg
                      Reply  Tue Apr 17 09:57:09 2018, Steve, Update, General, seismometer interfaces BS_Tril_Intrf-1X5.jpgGurs_Intf-1X1.jpg
                         Reply  Wed Apr 18 20:33:19 2018, Kevin, Update, General, seismometer interfaces front.jpgback.jpgconnectors.jpg
                            Reply  Fri Apr 20 23:36:28 2018, Kevin, Update, PEM, Seismometer BLRMs front.jpgback.jpgdetail.jpg
                               Reply  Tue Apr 24 21:19:08 2018, Kevin, Update, PEM, Seismometer BLRMs modified_schematic.pdftop.jpgbottom.jpgBS_Seis_PSD.pdf
                                  Reply  Wed Apr 25 17:44:39 2018, Arnold, Update, PEM, PEM Anti-Alias wiring 
                                     Reply  Thu Apr 26 09:35:49 2018, Kevin, Update, PEM, PEM Anti-Alias wiring board.jpgSeismometerPSD.pdf
                                        Reply  Thu Apr 26 20:22:21 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections setup.pdfseismometerASD.pdf
                                           Reply  Wed May 16 21:02:22 2018, Kevin, Update, PEM, ADC common mode rejection with new seismometer connections ADC_noise.pdf
                                        Reply  Thu Jun 14 15:24:32 2018, Steve, Update, PEM, ADC DAC In Line Test Boards are in ADC_DAC_in_(1).JPG
    Reply  Tue Jan 2 15:43:35 2018, Steve, Update, VAC, pumpdown after power outage pumpdown_from_44_mTorr.png
       Reply  Wed Feb 7 15:50:42 2018, Steve, Update, VAC, IFO pressure monitoring CC_analog.png
Message ID: 13615     Entry time: Wed Feb 7 15:50:42 2018     In reply to: 13495
Author: Steve 
Type: Update 
Category: VAC 
Subject: IFO pressure monitoring 

Hornet cold cathode gauge analoge output  [ DSub9 pin 3 and 7 ] are wired to go ETMX Acromag. It was reading 4.9V at 7.8e-6 Torr [ 3,110 V  8.35e-5A ] at the end of a 24ft BNC cable. Now it has to be hook up to an Acromag channel.

This will replace the not functioning C1: Vac-CC1_pressure

gautam: the motivation behind hooking this gauge up to our DAQ system is that non-vacuum-system-experts have a quick diagnostic to make sure everything is in order. This gauge is physically placed adjacent to V1, and so if something goes wrong with our vacuum pumps, we would see the effect here immediately. we did note that occassionally, the reading fluctuated by ~1V on the DMM used to check the voltage output at the end of the BNC cable, so we still need to run some long-term stability analysis once this channel is hooked up to the Acromag.  For future reference, in order to make this gauge work, we need to check that

  1. Error flag has been cleared.
  2. HV is ON (state has to be manually toggled).
Quote:

 

Quote:

There was a power outage.

The IFO pressure is 12.8 mTorr-it and it is not pumped. V1 is still closed. TP1 is not running. The Rga is not powered.

The PSL output shutter is still closed. 2W Innolight turned on and manual beam block placed in its beampath.

3 AC units turned on at room temp 84F

IFO pumped down from 44 mTorr to 9.6e-6 Torr with Maglev  backed with only TP3

Aux drypump  was helping our std drypump during this 1 hour period. TP3 reached 32 C and slowed down 47K rpm

The peak foreline pressure at P2  was ~3 Torr

Hornet cold cathode gauge setting:   research mode, air,

                                                            2830 HV  1e-4A  at 9.6e-6 Torr,

                                                         [  3110 HV  8e-5A at 7.4e-6 Torr one day later ]

Annuloses are at 2 Torr, not pumped

Valve configuration:  vacuum normal, RGA is still off

PSL shutter is opened automatically. Manual block removed.

End IR lasers and doublers are turned on.

 

NOTE: Maglev " rotation X " on vacuum medm screen is not working! " C1:Vac-TP1_rot " channel was removed.  Use " NORMAL X " for rotation monitoring.

*We removed this (i.e. rotation) field from the MEDM screen to avoid confusion.

 

Attachment 1: CC_analog.png  145 kB  | Hide | Hide all
CC_analog.png
ELOG V3.1.3-